The coastline of California is 46 miles longer than twice the length of Louisiana's coastline. It is also 443 miles longer than Louisiana's coastline. Find the length of California and Louisiana.
Answer by checkley71(8403)

C=2L+46
C=L+443
SEING AS WE HAVE 2 EQUATIONS EQUAL TO C THEN WE SET THESE 2 EQUAL TO EACH OTHER.
2L+46=L+443
2L-L=443-46
L=397 IS THE LENGTH OF LOUISIANA'S COAST.
C=2*397+46=840 FOR CALIFORNIA'S COAST LINE.
PROOF
840=397+443
840=840